sin(x+15 degrees) = cos(2x+30 degrees)

Answer by lwsshak3(11628)

sin(x+15)degrees = cos(2x+30 degrees)
Identity: sinx=cos(90-x)
sin(x+15)degrees=cos(90-(x+15))deg=cos(2x+30) deg
90-x-15=2x+30
3x=45
x=15 deg